OpenStudy (anonymous):

What will be the value of an account be after 8 years if $100 is invested at 6.0% interest compound continuously.... Please help!! Thanks! :)

OpenStudy (anonymous):

FV = limit P(1 + 1/m)Ymr m>oo

OpenStudy (anonymous):

@mathstudent55 is this will be the same formula?

OpenStudy (mathstudent55):

Why don't you follow exactly what was done for the previous similar problem. Try it, and post here your steps, and I'll check it.

OpenStudy (anonymous):

A = P*e^(Yr) where A = amount P = prinicipal Y = number of years r = interest rat expressed as a decima; fraction e = known constant plugging in, A = 100*e^(8*0.06) = 161.61 $ ans: $ 161.61

OpenStudy (mathstudent55):

Wait, it's a similar problem in that it uses the same formula, but you are given different info and are looking for different info. Start with the same formula A = pe^(rt)

OpenStudy (mathstudent55):

p = 100 t = 8 r = 6% = 0.06 you are looking for A

OpenStudy (anonymous):

answer is $ 161.61, I am sure

OpenStudy (anonymous):

@blast234 yes it is, and the details are clear, Thanks!!

OpenStudy (anonymous):

ur welcome :D

OpenStudy (anonymous):

@mathstudent55 Thanks also for your help... :*

OpenStudy (anonymous):

I have more don't worry guys!!! lol!

OpenStudy (anonymous):

lol